### v1

# Data-Driven Exploration for a Class of Continuous-Time Linear–Quadratic Reinforcement Learning Problems

Abstract
We study reinforcement learning (RL) for the same class of continuous-time stochastic linear–quadratic (LQ) control problems as in [1], where volatilities depend on both states and controls while states are scalar-valued and running control rewards are absent. We propose a model-free, data-driven exploration mechanism that adaptively adjusts entropy regularization by the critic and policy variance by the actor. Unlike the constant or deterministic exploration schedules employed in [1], which require extensive tuning for implementations and ignore learning progresses during iterations, our adaptive exploratory approach boosts learning efficiency with minimal tuning. Despite its flexibility, our method achieves a sublinear regret bound that matches the best-known model-free results for this class of LQ problems, which were previously derived only with fixed exploration schedules. Numerical experiments demonstrate that adaptive explorations accelerate convergence and improve regret performance compared to the non-adaptive model-free and model-based counterparts.
Keywords: C ontinuous-Time Reinforcement Learning, Stochastic Linear–-Quadratic Control, Actor, Critic, Data-Driven Exploration, Model-Free Policy Gradient, Regret Bounds.
1 Introduction
Linear–quadratic (LQ) control is a cornerstone of optimal control theory, widely applied in engineering, economics, and robotics due to its analytical tractability and practical relevance. Its theory has been extensively developed in the classical model-based setting where all the LQ coefficients are known and given [2, 3]. However, in real-world applications, complete knowledge of model parameters is rarely available. Many environments exhibit LQ-like structure yet key parameters in dynamics and objectives may be partially known or entirely unknown. While one can use historical data to estimate those model parameters as commonly done in the so-called “plug-in” approach, estimating some of the parameters to a required accuracy may not be possible. For instance, it is statistically impossible to estimate a stock return rate for a very small period – a phenomenon termed “mean-blur” [4, 5]. On the other hand, objective function values can be highly sensitive to estimated model parameters, rendering inferior performances or even complete irrelevance of model-based solutions [6].
Given these challenges, model-free reinforcement learning (RL) has emerged as a promising alternative for stochastic control, which directly learns optimal policies through data without estimating model parameters. A centerpiece of RL is exploration, with which the agent interacts with the unknown environment and improves her policies. The simplest exploration strategy is perhaps the so-called $\epsilon$-greedy for bandits problem [7]. More sophisticated and efficient exploration strategies include intrinsic motivation methods that reward exploring novel or unpredictable states [8, 9, 10, 11, 12, 13, 14], count-based exploration that encourages the agent to explore less frequently visited states [15, 16], go-explore that explicitly stores and revisits promising past states to enhance sample efficiency [17, 18], imitation-based exploration which leverages expert demonstrations [19], and safe exploration methods that rely on predefined safety constraints or risk models [20, 21]. While these methods are shown to be successful in their respective contexts, they are all for discrete-time problems and seem difficult to be extended to the continuous-time counterparts. However, most real-life applications are inherently continuous-time with continuous state spaces and possibly continuous control spaces (e.g., autonomous driving, robot navigation, and video game playing). On the other hand, transforming a continuous-time problem into a discrete-time one upfront by discretization has shown to be problematic when the time step size becomes very small [22, 23, 24].
Thus, there is a pressing need for developing exploration strategies tailored to continuous-time RL, particularly and foremost in the LQ setting, for achieving both stability and learning efficiency. Recent works, such as [1, 25, 26], have employed constant or deterministic exploration schedules to regulate exploration parameters and proved to achieve sublinear regret bounds. However, these approaches come with notable drawbacks. They require extensive manual tuning of the exploration hyperparameters, which considerably increases computational costs, yet these tuning efforts are typically not reflected in theoretical analyses and results including those concerning regret bounds. Moreover, pre-determined exploration strategies lack adaptability by nature, as they do not react to the incoming data nor to the current learning states, typically resulting in slower convergence and increased variance from excessive exploration or premature convergence to suboptimal policies from insufficient exploration.
This paper presents a companion research of [1]. We continue to study the same class of stochastic LQ problems in [1], but develop a data-driven adaptive exploration framework for both critic (the value functions) and actor (the control policies). Specifically, we propose a data-driven adaptive exploration strategy that dynamically adjusts the entropy regularization parameter for the critic and the stochastic policy variance for the actor based on the agent’s learning progress. Moreover, we provide theoretical guarantees by proving the almost sure convergence of both the adaptive exploration and policy parameters, along with a sublinear regret bound of $O(N^{\frac{3}{4}})$. This bound matches that in [1] whose analysis relies on fixed/deterministic exploration schedules. Finally, we validate our theoretical results through numerical experiments, demonstrating a faster convergence and improved regret bounds compared with a model-based benchmark with fixed exploration. For a comparison with the previous paper [1], we design experiments with both fixed and randomly generated model parameters. While theoretically the regret bounds in the two papers are of the same order, the numerical experiments show that our adaptive exploration strategy consistently outperforms and achieves lower regrets.
The remainder of this paper is organized as follows. Section 2 formulates the problem. Section 3 discusses the limitations of deterministic exploration strategies and presents our data-driven adaptive exploration mechanism. Section 4 describes the continuous-time LQ-RL algorithm with adaptive exploration. Section 5 establishes theoretical guarantees, proving convergence properties and regret bounds. Section 6 presents numerical experiments that validate the effectiveness of our approach and compare it against model-free and model-based benchmarks with deterministic exploration schedules. Finally, Section 7 concludes.
2 Problem Formulation and Preliminaries
2.1 Stochastic LQ Control: Classical Setting
We consider the same stochastic LQ control problem in [1]. The one-dimensional state process $x^{u}=\{x^{u}(t)\in\mathbb{R}:0\leq t\leq T\}$ evolves under the multi-dimensional control process $u=\{u(t)\in\mathbb{R}^{l}:0\leq t\leq T\}$ according to the stochastic differential equation (SDE):
$\mathrm{d}x^{u}(t)=$ $(Ax^{u}(t)+B^{\top}u(t))\mathrm{d}t+\sum_{j=1}^{m}(C_{j}x^{u}(t)+D_{j}^{\top}u(t))\mathrm{d}W^{(j)}(t),$ (1)
where $A$ and $C_{j}$ are scalars, $B$ and $D_{j}$ are $l\times 1$ vectors, and the initial condition is $x^{u}(0)=x_{0}$. The noise process $W=\{(W^{(1)}(t),\dots,W^{(m)}(t))^{\top}\in\mathbb{R}^{m}:0\leq t\leq T\}$ is an $m$-dimensional standard Brownian motion. We assume $\sum_{j=1}^{m}D_{j}D_{j}^{\top}>0$ to ensure well-posedness of the problem soon to be formulated.
The objective is to find a control process $u$ that maximizes the expected quadratic reward:
$\max_{u}\mathbb{E}\left[\int_{0}^{T}-\frac{1}{2}Qx^{u}(t)^{2}\mathrm{d}t-\frac{1}{2}Hx^{u}(T)^{2}\right],$ (2)
where $Q\geq 0$ and $H\geq 0$ are scalar weighting parameters. If the model parameters $A$, $B$, $C_{j}$, $D_{j}$, $Q$, and $H$ are known, the problem has an explicit solution as in [3, Chapter 6].
2.2 Stochastic LQ Control: RL Setting
We now consider the model-free, RL version of the problem just formulated, following [27], without assuming a complete knowledge of the model parameters or relying on parameter estimation.
In this case, there is no Riccati equation to solve (because its coefficients are unknown) and no explicit solutions to compute as in [3, Chapter 6]. Instead, the RL agent randomizes controls and considers distribution-valued control processes of the form $\pi=\{\pi(\cdot,t)\in\mathcal{P}(\mathbb{R}^{l}):0\leq t\leq T\}$, where $\mathcal{P}(\mathbb{R}^{l})$ is the set of all probability density functions over $\mathbb{R}^{l}$. The agent employs the control $u=\{u(t)\in\mathbb{R}^{l}:0\leq t\leq T\}$, where $u(t)\in\mathbb{R}^{l}$ is sampled from $\pi(\cdot,t)$ at each $t$, to control the original state dynamic.
According to [27], the system dynamic under a randomized control $\pi$ follows the SDE:
$\mathrm{d}x^{\pi}(t)=\widetilde{b}(x^{\pi}(t),\pi(\cdot,t))\mathrm{d}t+\sum_{j=1}^{m}\widetilde{\sigma}_{j}(x^{\pi}(t),\pi(\cdot,t))\mathrm{d}W^{(j)}(t),$ (3)
where
$\widetilde{b}(x,\pi):=Ax+B^{\top}\int_{\mathbb{R}^{l}}u\pi(u)\mathrm{d}u,$ (4)
$\widetilde{\sigma}_{j}(x,\pi):=\sqrt{\int_{\mathbb{R}^{l}}(C_{j}x+D_{j}^{\top}u)^{2}\pi(u)\mathrm{d}u}.$ (5)
To encourage exploration, an entropy term is added to the objective function. The entropy-regularized value function associated with a given policy $\pi$ is defined as
$J(t,x;\pi)=\mathbb{E}\biggl{[}$ $\int_{t}^{T}\left(-\frac{1}{2}Q(x^{\pi}(s))^{2}+\gamma p^{\pi}(s)\right)\mathrm{d}s-\frac{1}{2}H(x^{\pi}(T))^{2}\Big{|}x^{\pi}(t)=x\biggl{]},$ (6)
where $p^{\pi}(t)=-\int_{\mathbb{R}^{l}}\pi(t,u)\log\pi(t,u)\mathrm{d}u$ denotes the differential entropy of $\pi$, and $\gamma\geq 0$ is the so-called temperature parameter, which represents the trade-off between exploration and exploitation. The optimal value function is then given by $V(t,x)=\max_{\pi}J(t,x;\pi).$
Applying standard stochastic control techniques, [1] derives explicitly the optimal value function and optimal stochastic feedback policy of the above problem:
$V(t,x)=-\frac{1}{2}k_{1}(t)x^{2}+k_{3}(t),$ (7)
$\pi(u\mid t,x)=\mathcal{N}\left(u\Big{|}\bar{\mu}x,\Sigma\right),$ (8)
where $\mathcal{N}(\cdot|\bar{\mu}x,\Sigma)$ represents a multivariate Gaussian distribution with mean $\bar{\mu}x$ and covariance matrix $\Sigma$ with $\bar{\mu}=-\left(\sum_{j=1}^{m}D_{j}D_{j}^{\top}\right)^{-1}\left(B+\sum_{j=1}^{m}C_{j}D_{j}\right)$ and $\Sigma=\frac{\gamma}{k_{1}(t)}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}$. The functions $k_{1}$ and $k_{3}$ are determined by ordinary differential equations:
$k_{1}^{\prime}(t)=$ $-\biggl{[}2A+2B^{\top}\bar{\mu}+\sum_{j=1}^{m}\biggl{(}D_{j}^{\top}\bar{\mu}\bar{\mu}^{\top}D_{j}+2C_{j}D_{j}^{\top}\bar{\mu}+C_{j}^{2}\biggr{)}\biggr{]}k_{1}(t)-Q,\qquad k_{1}(T)=H,$ (9)
$k_{3}^{\prime}(t)=$ $\frac{k_{1}(t)}{2}\sum_{j=1}^{m}D_{j}^{\top}\Sigma D_{j}-\frac{\gamma}{2}\log\biggl{(}(2\pi e)^{l}\det(\Sigma)\biggl{)},\qquad k_{3}(T)=0.$
It is clear from (9) that $k_{1}(t)>0$ is independent of $\gamma$, and $k_{1}$, $k_{3}$ along with their derivatives are uniformly bounded over $[0,T]$.
We stress that the above are theoretical results that cannot be used to compute the optimal solutions because none of the parameters in the expressions is known; yet they offer valuable structural insights for developing the LQ-RL algorithms. Specifically, the optimal value function (critic) is quadratic in the state variable $x$, while the optimal (randomized) feedback control policy (actor) follows a Gaussian distribution, with its mean exhibiting a linear relationship with $x$. This intrinsic structure will play a pivotal role in reducing the complexity of function parameterization/approximation in our subsequent learning procedure.
2.3 Exploration in LQ-RL: Actor and Critic Perspectives
In our RL formulation above, both the actor and the critic engage in controls of exploration. The actor does this through the variance in the Gaussian policy (8), which represents the level of additional randomness injected into the interactions with the environment. The critic, on the other hand, regulates exploration via the entropy-regularized objective (6), where the temperature parameter $\gamma$ determines, if indirectly, the extent of exploration. Managing exploration properly is both important and challenging: excessive exploration can cause instability and hinder convergence, and insufficient exploration may prevent effective policy improvement. The next section introduces a data-driven adaptive exploration strategy designed to address these challenges and optimize policy learning.
3 Data-driven Exploration Strategy for LQ-RL
This section presents the core contribution of this work: data-driven explorations by both the actor and the critic.
3.1 Limitations of Deterministic Exploration Strategies
A critical challenge in RL is to design exploration strategies that are both effective and efficient. In the setting of continuous-time LQ, existing methods [1, 25] employ non-adaptive strategies by setting exploration levels of both actors and critics either as fixed constants or some deterministic schedules of time. While these strategies are proved to have theoretical guarantees, they exhibit several essential limitations in learning efficiency and practical implementations.
One major limitation of deterministic explorations lies in their arbitrary nature. Whether fixed as a constant or a predefined sequence, such a schedule often requires extensive manual tuning in actual implementations. The tuning process in turn may introduce significant computational and time costs for learning; yet these costs are rarely incorporated into theoretical analyses, creating a disconnect between theoretical guarantees and practical performance.
Another drawback of deterministic explorations is their lack of adaptability, as they follow fixed exploration trajectories regardless of the current states of the iterates, often leading to both excessive and insufficient explorations. While maintaining a constant exploration level is clearly ineffective, a deterministic schedule also suffers from a fundamental limitation: it adjusts the exploration parameter in a predetermined direction. Consequently, when the current exploration level is either too high or too low, the fixed schedule not only fails to correct this but may indeed exacerbate the situation by moving in the wrong direction. We will illustrate this issue with the experiments presented in Section 6.3.
To address these limitations, we propose an endogenous, data-driven approach that adaptively updates both the actor and critic exploration parameters, which will be shown in the subsequent sections.
3.2 Adaptive Critic Exploration
Critic Parameterization
Motivated by the form of the optimal value function in (7), we parameterize the value function with learnable parameters $\boldsymbol{\theta}\in\mathbb{R}^{d}$ and temperature parameter $\gamma\in\mathbb{R}$:
${J}(t,x;\boldsymbol{\theta},\gamma)=-\frac{1}{2}{k}_{1}(t;\boldsymbol{\theta})x^{2}+{k}_{3}(t;\boldsymbol{\theta},\gamma),$ (10)
where both $k_{1}$ and $k_{3}$ and their derivatives are continuous. Both functions can be neural nets; e.g. $k_{1}(\cdot;\boldsymbol{\theta})$ can be, for example, a positive neural network with a sigmoid activation. Furthermore, these functions are constructed so that there are appropriate positive constants $c_{1},c_{2},c_{3}$ such that
$\frac{1}{c_{2}}\leq{k}_{1}(t;\boldsymbol{\theta})\leq{c}_{2},\big{|}{k}_{1}^{\prime}(t;\boldsymbol{\theta})\big{|}\leq{c}_{1},\big{|}{k}_{3}^{\prime}(t;\boldsymbol{\theta},\gamma)\big{|}\leq{c}_{3},$ (11)
for all $0\leq t\leq T$, $|\boldsymbol{\theta}|\leq c^{(\boldsymbol{\theta})}$, and $|\gamma|\leq c^{(\gamma)}$, where $c^{(\boldsymbol{\theta})}>0$ and $c^{(\gamma)}>0$ are (fixed) hyperparameters. The values of $c_{1},c_{2},c_{3}$ are determined by the specific parametric forms of $k_{1}(\cdot;\boldsymbol{\theta})$ and $k_{3}(\cdot;\boldsymbol{\theta},\gamma)$, as well as $c^{(\boldsymbol{\theta})}$ and $c^{(\gamma)}$.  [^1] The values of $c_{1},c_{2},c_{3},c^{(\boldsymbol{\theta})},c^{(\gamma)}$ do not affect the order of regret bound, which will be shown in the subsequent sections. The key consideration here is the boundedness of $k_{1}$, $k_{1}^{\prime}$ and $k_{3}^{\prime}$. The boundedness assumptions (11) follow from the fact that when the model parameters are known, the corresponding functions satisfy the same conditions, as shown in Section 2.2.
In what follows, we introduce the subscript $n$ to denote values at the $n$-th iteration. For instance, $\gamma_{n}$ represents the updated value of $\gamma$ at iteration $n$, which evolves dynamically as learning progresses.
Data-Driven Temperature Parameter
The temperature parameter $\gamma$ plays a crucial role in RL in controlling the weight on the entropy-regularized term in the objective function (6). It governs the level of exploration by the critic for the purpose of balancing between exploration and exploitation.
In the related literature, $\gamma$ has been either taken as an exogenous constant hyperparameter ([1, 28, 29]) or a deterministic sequence of $n$ ([25]), which in turn requires extensive tunings in implementations. By contrast, we derive an adaptive temperature schedule, which is endogenous and data-driven, adjusted dynamically over iterations. Specifically, the adaptive update of $\gamma_{n}$ is determined by the values of the random processes $\boldsymbol{\theta}_{n}$ and a predefined deterministic sequence $b_{n}$ as follows
$\gamma_{n}=\frac{c_{\gamma}\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathrm{d}t}{b_{n}T},\quad\text{for }n=0,1,2,\dots.$ (12)
where $c_{\gamma}$ is a sufficiently large constant ensuring $\frac{1}{c_{\gamma}}$ is smaller than the minimum eigenvalue of $(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}$, and $b_{n}$ is a monotonically increasing sequence satisfying $b_{n}\uparrow\infty$ to be specified shortly. As will become clear in the subsequent sections, this updating rule is chosen to ensure convergence and achieve the desired sublinear regret bounds.
Incidentally, it follows from (11) that
$\gamma_{n}=\frac{c_{\gamma}\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\,\mathrm{d}t}{b_{n}T}\leq\frac{c_{\gamma}c_{2}}{b_{n}}\rightarrow 0$ (13)
as $n\rightarrow+\infty$.
3.3 Adaptive Actor Exploration
Actor Parameterization
Motivated by the structure of the optimal policy in (8), we parameterize the actor using learnable parameters $\boldsymbol{\phi}\in\mathbb{R}^{l}$ and $\Gamma\in\mathbb{S}^{l}_{++}$:
${\pi}(u\mid x;\boldsymbol{\phi},\Gamma)=\mathcal{N}(u\mid\boldsymbol{\phi}x,\Gamma).$ (14)
We write the corresponding entropy as $p^{\pi}(t)=p(t;\boldsymbol{\phi},\Gamma)$.
Adaptive Actor Exploration
The variance parameter $\Gamma$ represents the level of exploration controlled by the actor. A higher variance encourages broader exploration, allowing the agent to sample diverse actions, while a lower variance focuses more on exploitation. In the existing literature ([1, 25]), $\Gamma_{n}$ is chosen to be a deterministic sequence of $n$. Such a predetermined schedule does not account for the agent’s experience and observed data. Introducing an adaptive $\Gamma_{n}$ enables a more flexible and efficient exploration strategy.
To achieve this, we employ the general policy gradient approach developed in [30], which provides a foundation for updating $\Gamma$ in a fully data-driven manner. It leads to the following moment condition:
$\mathbb{E}$ $\biggl{[}\int_{0}^{T}\biggl{\{}\frac{\partial\log\pi}{\partial\Gamma}\left(u(t)\mid x(t);\boldsymbol{\phi},\Gamma\right)(\mathrm{d}J(t,x(t);\boldsymbol{\theta},\gamma)-\frac{1}{2}Qx(t)^{2}\mathrm{d}t+\gamma p(t;\boldsymbol{\phi},\Gamma)\mathrm{d}t)+\gamma\frac{\partial p}{\partial\Gamma}(t;\boldsymbol{\phi},\Gamma)\mathrm{d}t\biggr{\}}\biggr{]}=0.$ (15)
To enhance numerical efficiency in the resulting stochastic approximation algorithm, we reparametrize $\Gamma$ as $\Gamma^{-1}$. The chain rule implies that the derivative of $\Gamma^{-1}$ with respect to $\Gamma$ is a deterministic and time-invariant term, which can thus be omitted while preserving the validity of (15). Consequently,
[…]
$-\frac{1}{2}Qx(t)^{2}\mathrm{d}t+\gamma p(t;\boldsymbol{\phi},\Gamma)\mathrm{d}t)+\gamma\frac{\partial p}{\partial\Gamma^{-1}}(t;\boldsymbol{\phi},\Gamma)\mathrm{d}t\biggr{\}}\biggr{]}=0.$
The corresponding stochastic approximation algorithm then gives rise to the updating rule for the actor exploration parameter $\Gamma$:
$\Gamma_{n+1}\leftarrow\Gamma_{n}-a^{(\Gamma)}_{n}Z_{n}(T),$ (16)
where $a^{(\Gamma)}_{n}$ is the learning rate, and
[…]
$+\gamma_{n}p\left(t;\boldsymbol{\phi}_{n},\Gamma_{n}\right)\mathrm{d}t\biggr{]}+\gamma_{n}\frac{\partial p}{\partial\Gamma^{-1}}\left(t;\boldsymbol{\phi}_{n},\Gamma_{n}\right)\mathrm{d}t\biggl{\}},$
where $0\leq s\leq T$.
In sharp contrast with [1], the fact that $\Gamma_{n}$ is no longer a deterministic sequence that monotonically decreases to zero introduces significant analytical challenges. Specifically, the learning rates must be carefully chosen, and the convergence and convergence rate of $\Gamma_{n}$ must be carefully analyzed to ensure the algorithm finally still achieves a sublinear regret bound. These problems, along with their implications for overall performance, will be discussed in the subsequent sections.
4 A Continuous-Time RL Algorithm with Data-driven Exploration
This section presents the development of continuous-time RL algorithms with data-driven exploration. We discuss the policy evaluation and policy improvement steps, followed by a projection technique, time discretization, and the final pseudocode.
4.1 Policy Evaluation
To update the critic parameter $\boldsymbol{\theta}$, we employ policy evaluation (PE), a fundamental component of RL that focuses on learning the value function of a given policy.
Based on the parameterizations of the value function and policy in (10) and (14) respectively, we update $\boldsymbol{\theta}$ by applying the general continuous-time PE algorithm introduced by [31] along with stochastic approximation:
$\boldsymbol{\theta}_{n+1}$ $\leftarrow\boldsymbol{\theta}_{n}+a^{(\boldsymbol{\theta})}_{n}\int_{0}^{T}\frac{\partial J}{\partial\boldsymbol{\theta}}(t,x_{n}(t);\boldsymbol{\theta}_{n},\gamma_{n})\biggl{[}-\frac{1}{2}Qx_{n}(t)^{2}\mathrm{d}t+\mathrm{d}J(t,x_{n}(t);\boldsymbol{\theta}_{n},\gamma_{n})-\gamma_{n}{p}(t;\boldsymbol{\phi}_{n},\Gamma_{n})\mathrm{d}t\biggr{]},$ (18)
where $a^{(\boldsymbol{\theta})}_{n}$ is the corresponding learning rate.
4.2 Policy Improvement
The remaining learnable parameter is the mean of the stochastic Gaussian policy, $\boldsymbol{\phi}$. We employ policy gradient (PG) established in [30] and utilize stochastic approximation to generate the update rule:
$\boldsymbol{\phi}\leftarrow\boldsymbol{\phi}+a^{(\boldsymbol{\phi})}_{n}Y_{n}(T),$ (19)
where $a^{(\boldsymbol{\phi})}_{n}$ is the learning rate and
[…]
$+\mathrm{d}J\left(t,x_{n}(t);\boldsymbol{\theta}_{n},\gamma_{n}\right)+\gamma_{n}p\left(t;\boldsymbol{\phi}_{n},\Gamma_{n}\right)\mathrm{d}t\biggr{]}+\gamma_{n}\frac{\partial p}{\partial\boldsymbol{\phi}}\left(t;\boldsymbol{\phi}_{n},\Gamma_{n}\right)\mathrm{d}t\biggl{\}},$
where $0\leq s\leq T$.
4.3 Projections
The parameter updates for $\boldsymbol{\theta}$, $\boldsymbol{\phi}$, and $\Gamma$ follow stochastic approximation (SA), a foundational method introduced by [32] and extensively studied in subsequent works [33, 34, 35]. However, directly applying SA in our setting presents challenges such as extreme state values and unbounded estimation errors, which can destabilize the learning process. To address these issues, we incorporate the projection technique proposed by [36], ensuring that parameter updates remain in a bounded region at every iteration.
Define $\Pi_{K}(x):=\arg\min_{y\in K}|y-x|^{2}$, the projection of a point $x$ onto a set $K$. We now introduce the projection sets for all the critic and actor parameters.
First, we define the projection sets for the critic parameter $\boldsymbol{\theta}$ and the temperature parameter $\gamma$:
$K^{(\boldsymbol{\theta})}=\left\{\boldsymbol{\theta}\in\mathbb{R}^{d}\mid|\boldsymbol{\theta}|\leq c^{(\boldsymbol{\theta})}\right\},K^{(\gamma)}=\left\{\gamma\in\mathbb{R}\mid|\gamma|\leq c^{(\gamma)}\right\},$ (21)
which are employed to enforce the boundedness conditions in (11). The update rules incorporating projection for $\boldsymbol{\theta}$ and $\gamma$ are modified to
$\begin{split}\boldsymbol{\theta}_{n+1}\leftarrow\Pi_{K^{(\boldsymbol{\theta})}}\biggl{(}&\boldsymbol{\theta}_{n}+a^{(\boldsymbol{\theta})}_{n}\int_{0}^{T}\frac{\partial J}{\partial\boldsymbol{\theta}}(t,x_{n}(t);\boldsymbol{\theta}_{n},\gamma_{n})\\ &\biggl{[}\mathrm{d}J\left(t,x_{n}(t);\boldsymbol{\theta}_{n},\gamma_{n}\right)-\frac{1}{2}Qx_{n}(t)^{2}\mathrm{d}t+\gamma_{n}{p}\left(t;\boldsymbol{\phi}_{n},\gamma_{n}\right)\mathrm{d}t\biggr{]}\biggl{)},\end{split}$ (22)
$\begin{split}\gamma_{n+1}\leftarrow\Pi_{K^{(\gamma)}}\biggl{(}\frac{c_{\gamma}\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathrm{d}t}{b_{n}T}\biggl{)}.\end{split}$ (23)
Next, we define the projection sets for the actor parameters $\boldsymbol{\phi}$ and $\Gamma$:
$K^{(\boldsymbol{\phi})}_{n}=\left\{\boldsymbol{\phi}_{n}\in\mathbb{R}^{l}\Big{|}|\boldsymbol{\phi}_{n}|\leq c^{(\boldsymbol{\phi})}_{n}\right\},$ (24)
$K^{(\Gamma)}_{n}=\left\{\Gamma_{n}\in\mathbb{S}^{l}_{++}\Big{|}|\Gamma_{n}|\leq c^{(\Gamma)}_{n},\Gamma_{n}-\frac{1}{b_{n}}I\in\mathbb{S}^{l}_{++}\right\},$
where $\{c^{(\boldsymbol{\phi})}_{n}\uparrow\infty\}$, $\{c^{(\Gamma)}_{n}\uparrow\infty\}$, and $\{b_{n}\uparrow\infty\}$ are positive, monotonically increasing sequences, with $b_{n}$ the same deterministic sequence in (12). The specifics of these sequences are given in Theorem 5.1. Clearly, the sequences of sets $K^{(\boldsymbol{\phi})}_{n}$ and $K^{(\Gamma)}_{n}$ expand over time, eventually covering the entire spaces $\mathbb{R}^{l}$ and $\mathbb{S}^{l}_{++}$, respectively. This ensures that our algorithm remains model-free without needing to have prior knowledge of the model parameters. The update rules with projection for $\boldsymbol{\phi}$ and $\Gamma$ are now
$\begin{split}\boldsymbol{\phi}_{n+1}\leftarrow&\Pi_{K^{(\boldsymbol{\phi})}_{n+1}}\biggl{(}\boldsymbol{\phi}_{n}+a^{(\boldsymbol{\phi})}_{n}Y_{n}(T)\biggl{)},\end{split}$ (25)
$\begin{split}\Gamma_{n+1}\leftarrow&\Pi_{K^{(\Gamma)}_{n+1}}\biggl{(}\Gamma_{n}-a^{(\Gamma)}_{n}Z_{n}(T)\biggl{)},\end{split}$ (26)
where $Y_{n}(T)$ and $Z_{n}(T)$ are respectively determined by (20) and (17).
4.4 Discretization
While our RL framework is continuous-time in both formulation and analysis, numerical implementation requires discretization at the final stage. To facilitate this, we divide the time horizon $[0,T]$ into uniform intervals of size $\Delta t_{n}$ at the $n$-th iteration, resulting in discretized rules: [^2] Analysis on discretization errors is covered and discussed in Theorems 5.3 - 5.8, and Remark 5.10.
$\boldsymbol{\theta}_{n+1}\leftarrow\Pi_{K^{(\boldsymbol{\theta})}}\biggl{(}$ $\boldsymbol{\theta}_{n}+a^{(\boldsymbol{\theta})}_{n}\sum_{k=0}^{\left\lfloor\frac{T}{\Delta t_{n}}-1\right\rfloor}\frac{\partial J}{\partial\boldsymbol{\theta}}(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n},\gamma_{n})\biggl{[}J\left(t_{k+1},x_{n}(t_{k+1});\boldsymbol{\theta}_{n},\gamma_{n}\right)-J\left(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n},\gamma_{n}\right)$ (27)
$-\frac{1}{2}Qx_{n}(t_{k})^{2}\Delta t_{n}+\gamma_{n}{p}\left(t_{k};\boldsymbol{\phi}_{n},\Gamma_{n}\right)\Delta t_{n}\biggl{]}\biggl{)},$
$\begin{split}\gamma_{n+1}\leftarrow\Pi_{K^{(\gamma)}}\biggl{(}\frac{c_{\gamma}}{b_{n}T}\sum_{k=0}^{\left\lfloor\frac{T}{\Delta t_{n}}-1\right\rfloor}k_{1}(t_{k};\boldsymbol{\theta}_{n})\Delta t_{n}\biggl{)}.\end{split}$ (28)
Moreover, $Y_{n}(T)$ and $Z_{n}(T)$ in the update rules (25) and (26) are approximated by
[…]
$\biggl{[}(J\left(t_{k+1},x_{n}(t_{k+1});\boldsymbol{\theta}_{n},\gamma_{n}\right)-J\left(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n},\gamma_{n}\right))\frac{1}{\Delta t_{n}}$
$-\frac{1}{2}Qx_{n}(t_{k})^{2}+\gamma_{n}{p}\left(t_{k};\boldsymbol{\phi}_{n},\Gamma_{n}\right)\biggl{]}+\gamma_{n}\frac{\partial{p}}{\partial\boldsymbol{\phi}}\left(t_{k};\boldsymbol{\phi}_{n},\Gamma_{n}\right)\biggl{\}},$
[…]
$\biggl{[}(J\left(t_{k+1},x_{n}(t_{k+1});\boldsymbol{\theta}_{n},\gamma_{n}\right)-J\left(t_{k},x_{n}(t_{k});\boldsymbol{\theta}_{n},\gamma_{n}\right))\frac{1}{\Delta t_{n}}$
$-\frac{1}{2}Qx_{n}(t_{k})^{2}+\gamma_{n}{p}\left(t_{k};\boldsymbol{\phi}_{n},\Gamma_{n}\right)\biggl{]}+\gamma_{n}\frac{\partial{p}}{\partial\Gamma^{-1}}\left(t_{k};\boldsymbol{\phi}_{n},\Gamma_{n}\right)\biggl{\}}.$
4.5 LQ-RL Algorithm Featuring Data-Driven Exploration
The preceding analysis gives rise to the following RL algorithm that integrates adaptive exploration:
Algorithm 1 Data-driven Exploration LQ-RL Algorithm
Input: Initial values of learnable parameters $\boldsymbol{\theta}_{0},\boldsymbol{\phi}_{0},\gamma_{0},\Gamma_{0}$
for $n=0$ to $N$ do
     Initialize $k=0$, time $t=t_{k}=0$, and state $x_{n}(t_{k})=x_{0}$.
     while $t<T$ do
         Sample action $u_{n}(t_{k})$ from policy by (Eq. (14)).
         Uupdate state $x_{n}(t_{k+1})$ by LQ dynamics (Eq. (1)).
         Update time: $t_{k+1}\leftarrow t_{k}+\Delta t$, and set $t\leftarrow t_{k+1}$.
     end while
     Collect trajectory data: $\{(t_{k},x_{n}(t_{k}),u_{n}(t_{k}))\}_{k\geq 0}$.
     Update value function parameters $\boldsymbol{\theta}_{n+1}$ by (Eq. (27)).
     Update policy mean parameters $\boldsymbol{\phi}_{n+1}$ by (Eq. (29)).
     Update critic exploration parameter $\gamma_{n+1}$ by (Eq. (28)).
     Update actor exploration parameter $\Gamma_{n+1}$ by (Eq. (30)).
end for
Output: Final parameters $\boldsymbol{\theta}_{N},\boldsymbol{\phi}_{N},\gamma_{N},\Gamma_{N}$.
5 Regret Analysis
This section contains the main contribution of this work: establishing a sublinear regret bound for the LQ-RL algorithm with data-driven exploration, presented in Algorithm 1. We prove that the algorithm achieves a regret bound of $O(N^{\frac{3}{4}})$ (up to a logarithmic factor), matching the best-known model-free bound for this class of continuous-time LQ problems as reported in [1].
To quickly recap the major differences from [1]: 1) the actor exploration parameter $\Gamma_{n}$ therein is a deterministic monotonically decreasing sequence of $n$ while we apply policy gradient for updating $\Gamma_{n}$; 2) the static temperature parameter $\gamma$ is used in [1], whereas in our algorithm $\gamma$ is adaptively updated; 3) [1] only considers the case when the initial state is nonzero ($x_{0}\neq 0$), and our analysis removes this assumption.
In the remainder of the paper, we use $c$ (and its variants) to denote generic positive constants. These constants depend solely on the model parameters $A,B,C_{j},D_{j},Q,H,x_{0},T,\gamma,m,l$, and their values may vary from one instance to another.
5.1 Convergence Analysis on $\gamma_{n}$ and $\Gamma_{n}$
To start, we explore the almost sure convergence of the critic and actor exploration parameters $\gamma_{n}$ and $\Gamma_{n}$, together with the convergence rate of $\Gamma_{n}$ in terms of the mean-squared error (MSE).
Theorem 5.1.
Consider Algorithm 1 with fixed positive hyperparameters $c_{1},c_{2},c_{3}$ and a sufficiently large fixed constant $c_{\gamma}$. Define the sequences:
$a^{(\Gamma)}_{n}=a^{(\boldsymbol{\phi})}_{n}=\frac{\alpha^{\frac{3}{4}}}{(n+\beta)^{\frac{3}{4}}},\quad b_{n}=1\vee\frac{(n+\beta)^{\frac{1}{4}}}{\alpha^{\frac{1}{4}}},$
$c^{(\boldsymbol{\phi})}_{n}=1\vee(\log\log n)^{\frac{1}{6}},c^{(\Gamma)}_{n}=1\vee\log n,\Delta t_{n}=T(n+1)^{-\frac{5}{8}},$
where $\alpha>0$ and $\beta>0$ are constants. Then, the following results hold:
  1. (a)
    As $n\to\infty$, $\gamma_{n}$ converges almost surely to 00, and $\Gamma_{n}$ converges almost surely to the zero matrix $\boldsymbol{0}$.
  2. (b)
    For all $n$, $\mathbb{E}[|\Gamma_{n}|^{2}]\leq c\frac{(\log n)^{p_{2}}(\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}},$ where $c$ and $p_{2}$ are positive constants.
These results guarantee that the adaptive exploration parameters $\gamma_{n}$ and $\Gamma_{n}$ diminish almost surely, and the convergence rate of $\Gamma_{n}$ is essential for the proof of the regret bound. The remainder of this subsection is devoted to the proof of Theorem 5.1.
Define the conditional expectation of $Z_{n}(T)$ given the current parameter iterates as
$h^{(\Gamma)}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})=\mathbb{E}[Z_{n}(T)\mid\boldsymbol{\theta}_{n},\boldsymbol{\phi}_{n},\gamma_{n},\Gamma_{n}],$
as well as the deviation
$\xi^{(\Gamma)}_{n}=Z_{n}(T)-h^{(\Gamma)}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n}).$
The update rule (26) can be rewritten as
$\Gamma_{n+1}=\Pi_{K^{(\Gamma)}_{n+1}}\left(\Gamma_{n}-a^{(\Gamma)}_{n}[h^{(\Gamma)}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})+\xi^{(\Gamma)}_{n}]\right).$ (31)
For reader’s convenience, we break the proof of Theorem 5.1 into several steps, most of which adapt the general stochastic approximation methodology (e.g., [36] and [37]) to the specific setting here.
5.1.1 A Variance Upper Bound
The following result establishes an upper bound on the variance of the increment $Z_{n}(T)$.
Lemma 5.2.
There exists a constant $c>0$ depending solely on the model parameters such that
$\operatorname{Var}\left(Z_{n}(T)\Big{|}\boldsymbol{\theta}_{n},\boldsymbol{\phi}_{n},\gamma_{n},\Gamma_{n}\right)\leq c\left(1+|\boldsymbol{\phi}_{n}|^{8}+|\Gamma_{n}|^{8}+(\log b_{n})^{8}+\gamma_{n}^{8}\right)\exp{\{c|\boldsymbol{\phi}_{n}|^{4}\}}.$ (32)
Proof.
The proof is similar to that of [1, Lemma B.2] except that we need to account for the estimates on $\Gamma_{n}$; so we will be brief here. It follows from (17) together with Ito’s lemma (applied to $J\left(t,x_{n}(t);\boldsymbol{\theta}_{n},\gamma_{n}\right)$) that
$\mathrm{d}Z_{n}(t)\triangleq$ $Z_{n}^{(1)}(t)\mathrm{d}t+\sum_{j=1}^{m}Z_{n}^{(2,j)}(t)\mathrm{d}W_{n}^{(j)}(t),$ (33)
for which we have the following estimates
$\mathbb{E}[|Z_{n}^{(1)}(t)|^{2}|$ $\boldsymbol{\theta}_{n},\boldsymbol{\phi}_{n},\gamma_{n},\Gamma_{n},x_{n}(t)]\leq c(1+|\Gamma_{n}|^{4})(1+x_{n}(t)^{4}+|\boldsymbol{\phi}_{n}|^{4}x_{n}(t)^{4}+\gamma_{n}^{2}+\gamma_{n}^{2}(\log(\det(\Gamma_{n})))^{2}),$
$\mathbb{E}$ $[|Z_{n}^{(2,j)}(t)|^{2}|\boldsymbol{\theta}_{n},\boldsymbol{\phi}_{n},\gamma_{n},\Gamma_{n},x_{n}(t)]\leq c(1+|\Gamma_{n}|^{3})(1+x_{n}(t)^{4}+|\boldsymbol{\phi}_{n}|^{2}x_{n}(t)^{4}+|\Gamma_{n}|^{2}x_{n}(t)^{4}+|\boldsymbol{\phi}_{n}|^{2}|\Gamma_{n}|^{2}x_{n}(t)^{2}).$
Taking expectations in the above conditional on $x_{n}(t)$ and applying [1, Lemma B.1], we get
$\mathbb{E}[|Z_{n}^{(1)}(t)|^{2}+\sum_{j=1}^{m}|Z_{n}^{(2,j)}(t)|^{2}|\boldsymbol{\theta}_{n},\boldsymbol{\phi}_{n},\gamma_{n},\Gamma_{n}]\leq c\left(1+|\boldsymbol{\phi}_{n}|^{8}+|\Gamma_{n}|^{8}+(\log b_{n})^{8}+\gamma_{n}^{8}\right)\exp{\{c|\boldsymbol{\phi}_{n}|^{4}\}}.$ (34)
This establishes the desired inequality. ∎
5.1.2 A Mean Increment
We now analyze the mean increment $h^{(\Gamma)}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})$ in the updating rule (31). Taking integration and expectation in (33):
$h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})=\frac{1}{2}\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathrm{d}t\Gamma_{n}\Bigg{(}\sum_{j}D_{j}D_{j}^{\top}\Bigg{)}\Gamma_{n}-\frac{\gamma_{n}T}{2}\Gamma_{n}.$ (35)
Given that $h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})$ is quadratic in $\Gamma_{n}$, we apply (11) to establish the bound:
$\left|h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})\right|\leq c(1+|\Gamma_{n}|^{2}+\gamma_{n}^{2}).$ (36)
5.1.3 Almost Sure Convergence of $\gamma_{n}$ and $\Gamma_{n}$
We now prove Part (a) of Theorem 5.1. Indeed we will present and prove a more general result that involves a bias term $\beta^{(\Gamma)}_{n}=\mathbb{E}[\xi^{(\Gamma)}_{n}\mid\mathcal{G}_{n}]$, which captures implementation errors such as the discretization error (see Remark 5.10). When $\beta^{(\Gamma)}_{n}=\mathbf{0}$, the result simplifies to Theorem 5.1-(a).
Theorem 5.3.
Assume $\xi^{(\Gamma)}_{n}$ satisfies $\mathbb{E}\left[\xi^{(\Gamma)}_{n}\Big{|}\mathcal{G}_{n}\right]=\beta^{(\Gamma)}_{n}$ and
$\mathbb{E}\left[\left|\xi^{(\Gamma)}_{n}-\beta^{(\Gamma)}_{n}\right|^{2}\Big{|}\mathcal{G}_{n}\right]\leq c(1+|\boldsymbol{\phi}_{n}|^{8}+|\Gamma_{n}|^{8}+(\log b_{n})^{8}+\gamma_{n}^{8})\exp{\{c|\boldsymbol{\phi}_{n}|^{4}\}},$ (37)
where $c>0$ is a constant independent of $n$, and $\{\mathcal{G}_{n}\}$ is the filtration generated by $\{\boldsymbol{\theta}_{m},\boldsymbol{\phi}_{m},\gamma_{m},\Gamma_{m};m=0,1,2,...,n\}$. Moreover, assume
$(i)$ $0<a^{(\Gamma)}_{n}\leq 1\text{ }\forall n,\sum_{n}a^{(\Gamma)}_{n}=\infty,\sum_{n}a^{(\Gamma)}_{n}|\beta^{(\Gamma)}_{n}|<\infty;$ (38)
$(ii)$ $c^{(\boldsymbol{\phi})}_{n}\uparrow\infty,c^{(\Gamma)}_{n}\uparrow\infty,\text{ and for }0\leq q_{1},q_{2},q_{3},0\leq q_{4}\leq 4,$
$\sum(a^{(\Gamma)}_{n})^{2}(c^{(\Gamma)}_{n})^{q_{1}}(c^{(\boldsymbol{\phi})}_{n})^{q_{2}}(\log b_{n})^{q_{3}}e^{c(c^{(\boldsymbol{\phi})}_{n})^{q_{4}}}<\infty;$
$(iii)$ $b_{n}\geq 1,b_{n}\uparrow\infty,\sum\frac{a^{(\Gamma)}_{n}}{b_{n}}=\infty,\sum|\frac{1}{b_{n}}-\frac{1}{b_{n+1}}|<\infty.$
Then, $\gamma_{n}\xrightarrow{a.s.}0$ and $\Gamma_{n}\xrightarrow{a.s.}\boldsymbol{0}$.
Proof.
It is straightforward to see that $\gamma_{n}$ almost surely converges to 00. To prove the almost sure convergence of $\Gamma_{n}$, the key idea is to establish inductive upper bounds on $|\Gamma_{n}|^{2}$, namely to bound $|\Gamma_{n+1}|^{2}$ by a function of $|\Gamma_{n}|^{2}$.
Define the deviation term $U^{(\Gamma)}_{n}=\Gamma_{n}-\Gamma_{n}^{*}$, where $\Gamma_{n}^{*}$ is given by:
$\Gamma_{n}^{*}=\frac{\gamma_{n}T}{\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathrm{d}t}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}.$ (39)
By the adaptive temperature parameter $\gamma_{n}$ in (12), $\Gamma_{n}^{*}$ can be rewritten as:
$\Gamma_{n}^{*}=\frac{c_{\gamma}}{b_{n}}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1},$ (40)
where we recall $c_{\gamma}>0$ is such that $\frac{1}{c_{\gamma}}$ is smaller than the minimum eigenvalue of $(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}$, denoted as $\lambda_{\min}((\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1})$. Thus $\Gamma_{n}^{*}>\frac{1}{b_{n}}I$. Since $b_{n}\uparrow\infty$, $\Gamma_{n}^{*}$ decreases in $n$, leading to the positive matrix difference: $G_{n}=\Gamma_{n}^{*}-\Gamma_{n+1}^{*}>0.$
We now show the almost sure convergence of $U^{(\Gamma)}_{n}$ to 0. Consider sufficiently large $n$ such that $\Gamma_{n}^{*}\in K^{(\Gamma)}_{n+1}$. By applying the general projection inequality $|\Pi_{K}(y)-x|^{2}\leq|y-x|^{2}$, we have
$\mathbb{E}\left[|U^{(\Gamma)}_{n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]$
$\leq$ $\mathbb{E}\left[|U^{(\Gamma)}_{n}-a^{(\Gamma)}_{n}[h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})+\xi^{(\Gamma)}_{n}]+G_{n}|^{2}\Big{|}\mathcal{G}_{n}\right]$
$\leq$ $|U^{(\Gamma)}_{n}|^{2}-2a^{(\Gamma)}_{n}\langle U^{(\Gamma)}_{n},h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})\rangle+(1+|U^{(\Gamma)}_{n}|^{2})(a^{(\Gamma)}_{n}|\beta^{(\Gamma)}_{n}|+|G_{n}|)$
$+4(a^{(\Gamma)}_{n})^{2}\biggl{(}|h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})|^{2}+|\beta^{(\Gamma)}_{n}|^{2}+|\frac{G_{n}}{a^{(\Gamma)}_{n}}|^{2}+\mathbb{E}\left[\left|\xi^{(\Gamma)}_{n}-\beta^{(\Gamma)}_{n}\right|^{2}\Big{|}\mathcal{G}_{n}\right]\biggr{)}.$
Recall that $\frac{1}{b_{n}}|\leq|\Gamma_{n}|\leq c^{(\Gamma)}_{n}$ almost surely. By (21), (24), (36), and (37), we can further derive that
$\mathbb{E}\left[|U^{(\Gamma)}_{n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]\leq(1+\kappa^{(\Gamma)}_{n})|U^{(\Gamma)}_{n}|^{2}-\zeta^{(\Gamma)}_{n}+\eta^{(\Gamma)}_{n},$
where $\kappa^{(\Gamma)}_{n}=a^{(\Gamma)}_{n}|\beta^{(\Gamma)}_{n}|+|G_{n}|$,
$\zeta^{(\Gamma)}_{n}=2a^{(\Gamma)}_{n}\langle U^{(\Gamma)}_{n},h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})\rangle$, and
[…]
$+|\beta^{(\Gamma)}_{n}|^{2}+|\frac{G_{n}}{a^{(\Gamma)}_{n}}|^{2}\biggl{\}}.$
Next, we prove that $\sum|G_{n}|<\infty$ and $\sum|G_{n}|^{2}<\infty$. By the assumption (iii) in (38) along with (40), we obtain
$\sum_{n=0}^{\infty}|G_{n}|$ $=\sum_{n=1}^{\infty}|\Gamma_{n}^{*}-\Gamma_{n+1}^{*}|<c\sum_{j=1}^{m}|\frac{1}{b_{n}}-\frac{1}{b_{n+1}}|<\infty.$ (41)
Furthermore, since $b_{n}\uparrow\infty$, $n_{0}=\inf{\{n^{\prime}\in\mathbb{N}:|G_{n}|<1\text{ for all }n\geq n^{\prime}\}}<\infty$. Therefore, (41) yields
$\sum_{n=1}^{\infty}|G_{n}|^{2}<\sum_{n=1}^{n_{0}-1}|G_{n}|^{2}+\sum_{n=n_{0}}^{\infty}|G_{n}|<\infty.$ (42)
By the assumptions (i)-(ii) of (38), as well as (41) and (42), we know $\sum\kappa^{(\Gamma)}_{n}<\infty$ and $\sum\eta^{(\Gamma)}_{n}<\infty$. It then follows from [38, Theorem 1] that $\left|U^{(\Gamma)}_{n}\right|^{2}$ converges to a finite limit and $\sum\zeta^{(\Gamma)}_{n}<\infty$ almost surely.
It suffices to show $|U^{(\Gamma)}_{n}|\to 0$ almost surely. The equations (24) and (LABEL:eq_h2_app) imply
$\zeta^{(\Gamma)}_{n}\geq$ $\frac{a^{(\Gamma)}_{n}T}{b_{n}c_{2}}\lambda_{\min}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})|U^{(\Gamma)}_{n}|^{2}.$
Now, suppose $|U^{(\Gamma)}_{n}|^{2}\rightarrow c$ almost surely, where $0<c<\infty$ is a constant. Then there exists a set $Z\in\mathcal{F}$ with $\mathbb{P}(Z)>0$ so that for every $\omega\in Z$, there exists a measurable set $Z\in\mathcal{F}$ with $\mathbb{P}(Z)=1$ such that, for every $\omega\in Z$, there exists a constant $0<\delta(\omega)<c$ satisfying
$|U^{(\Gamma)}_{n}|^{2}=|\Gamma_{n}-\Gamma_{n}^{*}|^{2}\geq c-\delta(\omega)>0\quad\text{for sufficiently large }n.$
Thus, by the assumption (iii) of (38),
$\sum\zeta^{(\Gamma)}_{n}\geq\frac{T}{c_{2}}(c-\delta(\omega))\lambda_{\min}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\sum\frac{a^{(\Gamma)}_{n}}{b_{n}}=\infty.$
This is a contradiction, proving that $\Gamma_{n}\xrightarrow{a.s.}\Gamma_{n}^{*}$. However, $\Gamma_{n}^{*}$ converges to $\boldsymbol{0}$ almost surely. Hence, $\Gamma_{n}\xrightarrow{a.s.}\boldsymbol{0}$ as well.
Remark 5.4.
A case satisfying the assumptions in (38) is when $a^{(\Gamma)}_{n}=1\wedge\frac{\alpha^{\frac{3}{4}}}{(n+\beta)^{\frac{3}{4}}}$ and $b_{n}=1\vee\frac{(n+\beta)^{\frac{1}{4}}}{\alpha^{\frac{1}{4}}}$, where constants $\alpha>0$, $\beta>0$. $c^{(\boldsymbol{\phi})}_{n}=1\vee(\log\log n)^{\frac{1}{6}},c^{(\Gamma)}_{n}=1\vee\log n,\beta^{(\boldsymbol{\phi})}_{n}=0$. This is because $\sum\frac{1}{n}=\infty$; $\sum\frac{(\log n)^{p}(\log\log n)^{q}}{n^{r}}<\infty$, for any $p,q>0$ and $r>1$; and $\sum|n^{-\frac{1}{4}}-(n+1)^{-\frac{1}{4}}|<\frac{1}{4}\sum n^{-\frac{5}{4}}<\infty$ (by the mean–value theorem). Additionally, it is interesting to note that the assumptions $\sum\frac{a^{(\Gamma)}_{n}}{b_{n}}=\infty$ and $\sum|\frac{1}{b_{n}}-\frac{1}{b_{n+1}}|<\infty$ are new compared to [1]. This is due to the implementation of data-driven exploration.
5.1.4 Convergence Rate of $\Gamma_{n}$
We need the following lemmas.
Lemma 5.5.
For any $W>0$ and any $10<q<$ 1, there exist positive numbers $\alpha>\frac{1}{W}$ and $\beta\geq\max(\frac{1}{W\alpha-1},W^{\frac{2q-1}{1-q}}\alpha^{\frac{q}{1-q}})$ such that the sequences $\hat{a}_{n}=\frac{\alpha}{n+\beta}$ and $a_{n}=\frac{\alpha^{q}}{(n+\beta)^{q}}$ satisfy $\hat{a}_{n}\leq\hat{a}_{n+1}(1+W\hat{a}_{n+1})$ and $a_{n}\leq a_{n+1}(1+Wa_{n+1})$ for all $n\geq 0$.
Proof.
First,
$\hat{a}_{n}\leq\hat{a}_{n+1}(1+W\hat{a}_{n+1})\Leftrightarrow n+1+\beta\leq W\alpha n+W\alpha\beta,$
the latter being true for any $n$ when $\alpha>\frac{1}{W}>0$, $\beta\geq\frac{1}{W\alpha-1}>0$.
Next, we have
$a_{n}\leq a_{n+1}(1+Wa_{n+1})$ (43)
$\Leftrightarrow$ $(n+\beta+1)^{q}-(n+\beta)^{q}\leq W\alpha^{q}\biggl{(}\frac{n+\beta}{n+\beta+1}\biggl{)}^{q}.$
For the latter inequality in (43), notice that the left-hand side decreases in $n$ while the right-hand side increases in $n$. So to show that this inequality is true for all $n$, it is sufficient to show that it is true when $n=0$, which is $(\beta+1)^{q}-\beta^{q}\leq W\alpha^{q}\frac{\beta^{q}}{(\beta+1)^{q}}$.
Now, for $\beta\geq\frac{1}{W\alpha-1}$, we have $\beta+1\leq W\alpha\beta$. On the other hand, the mean–value theorem yields $(\beta+1)^{q}-\beta^{q}\leq q\beta^{q-1}$. Hence,
$W^{\frac{2q-1}{1-q}}\alpha^{\frac{q}{1-q}}\leq$ $\beta\Rightarrow(\beta+1)^{q}-\beta^{q}\leq W\alpha^{q}\frac{\beta^{q}}{(\beta+1)^{q}}.$
Lemma 5.6.
Let the sequence $\hat{G}_{n}=n^{-q}-(n+1)^{-q}>0$ for some $q>0$. Then $\hat{G}_{n}^{2}=O(n^{-2q-2})$.
Proof.
We have
$\hat{G}_{n}^{2}=\left[\frac{(n+1)^{q}-n^{q}}{(n+1)^{q}n^{q}}\right]^{2}<\left[\frac{(n+1)^{q}-n^{q}}{n^{2q}}\right]^{2}.$
Applying the mean–value theorem to the function $f(x)=x^{q}$ and noting that $f^{\prime}(x)=qx^{q-1}$ is a decreasing function for $x>0$, we conclude
$\left[\frac{(n+1)^{q}-n^{q}}{n^{2q}}\right]^{2}<\frac{(qn^{q-1})^{2}}{n^{4q}}=q^{2}n^{-2q-2}.$
The following theorem specializes to Theorem 5.1-(b) when the bias term $\beta^{(\Gamma)}_{n}=0$.
Theorem 5.7.
Under the same setting of Theorem 5.3, if the sequence $\{\hat{a}_{n}\}$, defined as $\{\frac{a^{(\Gamma)}_{n}}{b_{n}}\}$, satisfies
$\hat{a}_{n}\leq\hat{a}_{n+1}(1+w\hat{a}_{n+1})$
for some sufficiently small constant $w>0$, and the sequences $\{\frac{b_{n}^{3}}{a^{(\Gamma)}_{n}}|\beta^{(\Gamma)}_{n}|^{2}\}$ and $\{\frac{b_{n}^{3}|G_{n}|^{2}}{(a^{(\Gamma)}_{n})^{3}}\}$ are non-decreasing in $n$, then there exists an increasing sequence $\{\hat{\eta}^{(\Gamma)}_{n}\}$ and a constant $c^{\prime}>0$ such that
$\mathbb{E}[|\Gamma_{n+1}-\Gamma_{n+1}^{*}|^{2}]\leq c^{\prime}\hat{a}_{n}\hat{\eta}^{(\Gamma)}_{n}.$
In particular, if the parameters $b_{n},a^{(\Gamma)}_{n},c^{(\boldsymbol{\phi})}_{n},c^{(\Gamma)}_{n},\beta^{(\Gamma)}_{n}$ are set as in Remark 5.4, then
$\mathbb{E}[|\Gamma_{n+1}|^{2}]\leq c\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}$
where $p_{2}$ is the same constant appearing in Theorem 5.3.
Proof.
First, it follows from (24) and (LABEL:eq_h2_app) that
$\langle U^{(\Gamma)}_{n},h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})\rangle\geq\frac{T}{2b_{n}c_{2}}\lambda_{\min}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})|U^{(\Gamma)}_{n}|^{2}:=\frac{\tilde{c}}{b_{n}}|U^{(\Gamma)}_{n}|^{2},$ (44)
where $\tilde{c}=\frac{T}{2c_{2}}\lambda_{\min}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})>0$.
Define $n_{1}=\inf\{n\in\mathbb{N}:\Gamma_{n}^{*}\in K^{(\Gamma)}_{n+1}\}$. For $n\geq n_{1}$, we follow a similar reasoning as in the proof of Theorem 5.3 to get
$\mathbb{E}\left[|U^{(\Gamma)}_{n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]\leq(1-\tilde{c}\frac{a^{(\Gamma)}_{n}}{b_{n}})|U^{(\Gamma)}_{n}|^{2}+4\frac{(a^{(\Gamma)}_{n})^{2}}{b_{n}^{2}}b_{n}^{2}\biggl{(}$ $|h^{(\Gamma)}(\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})|^{2}+(1+|\frac{b_{n}}{2\tilde{c}a^{(\Gamma)}_{n}}|)|\beta^{(\Gamma)}_{n}|^{2}$ (45)
$+|\frac{b_{n}G_{n}^{2}}{2\tilde{c}(a^{(\Gamma)}_{n})^{3}}|+|\frac{G_{n}}{a^{(\Gamma)}_{n}}|^{2}+\mathbb{E}\left[\left|\xi^{(\Gamma)}_{n+1}-\beta^{(\Gamma)}_{n}\right|^{2}\Big{|}\mathcal{G}_{n}\right]\biggl{)}.$
Moreover, the assumptions in (38) imply that $(1+|\frac{b_{n}}{2\tilde{c}a^{(\Gamma)}_{n}}|)|\beta^{(\Gamma)}_{n}|^{2}\leq c(\frac{b_{n}}{a^{(\Gamma)}_{n}}|\beta^{(\Gamma)}_{n}|^{2})$ and $\frac{|G_{n}|^{2}}{(a^{(\Gamma)}_{n})^{2}}\leq c(\frac{b_{n}|G_{n}|^{2}}{(a^{(\Gamma)}_{n})^{3}})$ for some constant $c>0$. When $n\geq n_{1}$, in view of (21), (24), (36), and (37), it follows from (45) that
$\mathbb{E}\left[|U^{(\Gamma)}_{n+1}|^{2}\Big{|}\mathcal{G}_{n}\right]\leq(1-\tilde{c}\hat{a}_{n})|U^{(\Gamma)}_{n}|^{2}+4\hat{a}_{n}^{2}\hat{\eta}^{(\Gamma)}_{n}$
, where
$\hat{\eta}^{(\Gamma)}_{n}=$ $cb_{n}^{2}\biggl{(}1+(c^{(\boldsymbol{\phi})}_{n})^{8}+(c^{(\Gamma)}_{n})^{8}+(\log b_{n})^{8}+\frac{b_{n}|\beta^{(\Gamma)}_{n}|^{2}}{a^{(\Gamma)}_{n}}+\frac{b_{n}|G_{n}|^{2}}{(a^{(\Gamma)}_{n})^{3}}\biggr{)}\exp{\{c(c^{(\boldsymbol{\phi})}_{n})^{4}\}},$ (46)
which is monotonically increasing because $c^{(\boldsymbol{\phi})}_{n}$, $c^{(\Gamma)}_{n}$, $b_{n}$ are monotonically increasing and $\frac{b_{n}^{3}}{a^{(\Gamma)}_{n}}|\beta^{(\Gamma)}_{n}|^{2}$, $\frac{b_{n}^{3}|G_{n}|^{2}}{(a^{(\Gamma)}_{n})^{3}}$ are non-decreasing by the assumptions. Taking expectations on both sides of the above and denoting $\rho_{n}=\mathbb{E}[|U^{(\Gamma)}_{n}|^{2}]$, we get
$\rho_{n+1}\leq(1-\tilde{c}\hat{a}_{n})\rho_{n}+4\hat{a}_{n}^{2}\hat{\eta}^{(\Gamma)}_{n}$ (47)
when $n\geq n_{1}$.
Next, we show $\rho_{n+1}\leq c^{\prime}\hat{a}_{n}\hat{\eta}^{(\Gamma)}_{n}$ for all $n\geq 0$, where $c^{\prime}=\max\{\frac{\rho_{1}}{\hat{a}_{0}\hat{\eta}^{(\Gamma)}_{0}},\frac{\rho_{2}}{\hat{a}_{1}\hat{\eta}^{(\Gamma)}_{1}},\cdots,\frac{\rho_{n_{1}+1}}{\hat{a}_{n_{1}}\hat{\eta}^{(\Gamma)}_{n_{1}}},\frac{4}{\tilde{c}}\}+1$. Indeed, it is true when $n\leq n_{1}$. Assume that $\rho_{k+1}\leq c^{\prime}a_{k}\hat{\eta}^{(\Gamma)}_{n}$ is true for $n_{1}\leq k\leq n-1$. Then (47) yields
$\rho_{n+1}$ $\leq(1-\tilde{c}\hat{a}_{n})\rho_{n}+4\hat{a}_{n}^{2}\hat{\eta}^{(\Gamma)}_{n}$
$\leq(1-\tilde{c}\hat{a}_{n})c^{\prime}\hat{a}_{n-1}\hat{\eta}_{1,n-1}+4\hat{a}_{n}^{2}\hat{\eta}^{(\Gamma)}_{n}$
$\leq(1-\tilde{c}\hat{a}_{n})c^{\prime}\hat{a}_{n}(1+w\hat{a}_{n})\hat{\eta}^{(\Gamma)}_{n}+4\hat{a}_{n}^{2}\hat{\eta}^{(\Gamma)}_{n}$
$=c^{\prime}\hat{a}_{n}\hat{\eta}^{(\Gamma)}_{n}+c^{\prime}\hat{\eta}^{(\Gamma)}_{n}\hat{a}_{n}^{2}\biggl{(}w-\tilde{c}-\tilde{c}w\hat{a}_{n}+\frac{4}{c^{\prime}}\biggl{)}.$
Consider the function
${f}(x)=c^{\prime}\hat{\eta}^{(\Gamma)}_{n}x^{2}\biggl{(}w-\tilde{c}-\tilde{c}wx+\frac{4}{c^{\prime}}\biggl{)},$
which has two roots at $x_{1,2}=0$ and one root at $x_{3}=\frac{w-(\tilde{c}-\frac{4}{c^{\prime}})}{cw}$. Because $\tilde{c}-\frac{4}{c^{\prime}}>0$, we can choose $0<w<\tilde{c}-\frac{4}{c^{\prime}}$ so that $x_{3}<0$. So ${f}(x)<0$ when $x>0$, leading to
$c^{\prime}\hat{\eta}^{(\Gamma)}_{n}\hat{a}_{n}^{2}\biggl{(}w-\tilde{c}-\tilde{c}w\hat{a}_{n}+\frac{4}{c^{\prime}}\biggl{)}<0,\;\;\forall n$
because $\hat{a}_{n}>0$. We have now proved $\mathbb{E}[|U^{(\boldsymbol{\phi})}_{n+1}|^{2}]\leq c^{\prime}\hat{a}_{n}\hat{\eta}^{(\Gamma)}_{n}$.
In particular, under the setting of Remark 5.4 and by Lemma 5.6, we can verify that $\frac{b_{n}^{3}}{a^{(\Gamma)}_{n}}|\beta^{(\Gamma)}_{n}|^{2}$ and $\frac{b_{n}^{3}|G_{n}|^{2}}{(a^{(\Gamma)}_{n})^{3}}$ are non-decreasing sequences of $n$, and $\hat{a}_{n}=\Theta(n^{-1})$. Then
$\hat{\eta}^{(\Gamma)}_{n}\leq cn^{\frac{1}{2}}(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}},$ (48)
where $c$ and $p_{2}$ are positive constants. Thus it follows that
$\mathbb{E}[|\Gamma_{n+1}-\Gamma_{n}^{*}|^{2}]\leq c\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}.$
Finally, noting that $\Gamma_{n}^{*}$ converges to $\boldsymbol{0}$ with the rate of $\frac{1}{b_{n}}$ by (40), we have
$\mathbb{E}[|\Gamma_{n}|^{2}]=$ $\mathbb{E}[|\Gamma_{n}-\Gamma_{n}^{*}+\Gamma_{n}^{*}|^{2}]\leq\mathbb{E}[|\Gamma_{n}-\Gamma_{n}^{*}|^{2}]+\mathbb{E}[|\Gamma_{n}^{*}|^{2}]$
$\leq$ $c^{\prime}\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}+c^{\prime}\frac{\alpha^{\frac{1}{2}}}{(n+\beta)^{\frac{1}{2}}}$
$\leq$ $c\frac{(e\vee\log n)^{p_{2}}(1\vee\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}}.$
The proof is complete.
5.2 Convergence Analysis of $\boldsymbol{\phi}_{n}$
In this sebsection, we investigate the convergence properties of the actor parameter $\boldsymbol{\phi}_{n}$ for both cases $x_{0}\neq 0$ and $x_{0}=0$.
Theorem 5.8.
Under the same setting of Theorem 5.1, we have
  1. (a)
    As $n\to\infty$, $\boldsymbol{\phi}_{n}$ converges almost surely to $\boldsymbol{\phi}^{*}=-\Bigg{(}\sum_{j=1}^{m}D_{j}D_{j}^{\top}\Bigg{)}^{-1}\Bigg{(}B+\sum_{j=1}^{m}C_{j}D_{j}\Bigg{)}.$
  2. (b)
    The expected squared error satisfies
    $\mathbb{E}[|\boldsymbol{\phi}_{n}-\boldsymbol{\phi}^{*}|^{2}]\leq c\frac{(\log n)^{p_{1}}(\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{2}}},\quad\text{if }x_{0}\neq 0,$
    $\mathbb{E}[|\boldsymbol{\phi}_{n}-\boldsymbol{\phi}^{*}|^{2}]\leq c\frac{(\log n)^{p_{1}^{\prime}}(\log\log n)^{\frac{4}{3}}}{n^{\frac{1}{4}}},\quad\text{if }x_{0}=0,$
    where $c$, $p_{1}$, and $p_{1}^{\prime}$ are positive constants only dependent on the model parameters.
Proof.
When $x_{0}\neq 0$, the proof is similar to that of [1, Theorem 4.1] with only minor modifications needed. When $x_{0}=0$, the proof becomes more delicate as the convergence behavior of $\boldsymbol{\phi}$ now depends explicitly on the actor exploration parameter $\Gamma$. Here, we highlight the differences needed.
Denoted
$h^{(\boldsymbol{\phi})}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})=\mathbb{E}[Y_{n}(T)\mid\boldsymbol{\theta}_{n},\boldsymbol{\phi}_{n},\gamma_{n},\Gamma_{n}],$
and the noise
$\xi^{(\boldsymbol{\phi})}_{n}=Y_{n}(T)-h^{(\boldsymbol{\phi})}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n}).$
The updating rule (25) for $\boldsymbol{\phi}$ is now
$\boldsymbol{\phi}_{n+1}=\Pi_{K^{(\boldsymbol{\phi})}_{n+1}}(\boldsymbol{\phi}_{n}+a^{(\boldsymbol{\phi})}_{n}[h^{(\boldsymbol{\phi})}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n}.\gamma_{n})+\xi^{(\boldsymbol{\phi})}_{n}]).$ (49)
Similar to [1, Section B], we obtain
$h^{(\boldsymbol{\phi})}(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})=-l(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})(\boldsymbol{\phi}_{n}-\boldsymbol{\phi}^{*}),$ (50)
where
$l(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})=(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\int_{0}^{T}k_{1}(t;\boldsymbol{\theta}_{n})\mathbb{E}[x_{n}(t)^{2}]\mathrm{d}t.$ (51)
When $x_{0}\neq 0$, we have $l(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})\geq\bar{c}I$, where $10<\bar{c}^{\prime}<$ 1 is independent of $\Gamma_{n}$. The same proof of [1, Theorem 4.1] then applies.
However, when $x_{0}=0$,
$l(\boldsymbol{\phi}_{n},\Gamma_{n};\boldsymbol{\theta}_{n},\gamma_{n})$ $\geq(\sum_{j=1}^{m}D_{j}D_{j}^{\top})\int_{0}^{T}\frac{c^{\prime}|\Gamma_{n}|t}{c_{2}}\mathrm{d}t=\frac{(\sum_{j=1}^{m}D_{j}D_{j}^{\top})T^{2}c^{\prime}}{2c_{2}}|\Gamma_{n}|\geq\bar{c}^{\prime}|\Gamma_{n}|I.$ (52)
Thus $l$ no longer admits a uniform lower bound strictly away from $\boldsymbol{0}$ because $\Gamma_{n}\rightarrow\boldsymbol{0}$. To get around, for establishing part (a), we make use of the condition $\sum\frac{a^{(\boldsymbol{\phi})}_{n}}{b_{n}}=\infty$, which ensures that the argument in proving [1, Theorem B.3] remains valid. (Notice that the hyperparameters $a^{(\boldsymbol{\phi})}_{n}$ and $b_{n}$ specified in Theorem 5.1 also satisfy this condition.) On the other hand, for proving part (b), we reinterpret $\hat{a}_{n}=\frac{a^{(\boldsymbol{\phi})}_{n}}{b_{n}}$ as the effective learning rate and follow the strategy outlined in the proof of Theorem 5.7. This adjustment allows the analysis in [1, Appendix B.6] to be carried over to our setting. ∎
The established convergence of the learned policy underpins the regret analysis of Algorithm 1. Theorem 5.8 posits that the MSE of $\boldsymbol{\phi}_{n}$ depends on whether the initial state $x_{0}$ is zero or nonzero; yet both cases ultimately lead to the same sublinear regret bound, which will be shown in the next subsection.
5.3 Regret Bound
The regret quantifies the (average) difference between the value function of the learned policy and that of the oracle optimal policy in the long run. For a stochastic Gaussian policy $\pi=\mathcal{N}(\cdot|\boldsymbol{\phi}x,\Gamma)$, we denote
$\bar{J}(\boldsymbol{\phi},\Gamma)=\mathbb{E}\biggl{[}$ $\int_{0}^{T}\left(-\frac{1}{2}Qx^{\pi}(s)^{2}\right)\mathrm{d}s-\frac{1}{2}Hx^{\pi}(T)^{2}\Big{|}x^{\pi}(0)=x_{0}\biggr{]}.$ (53)
This function evaluates policy $\mathcal{N}(\cdot|\boldsymbol{\phi}x,\Gamma)$ under the original objective without entropy regularization. The oracle value of the original problem is $\bar{J}(\boldsymbol{\phi}^{*},\boldsymbol{0})$.
Theorem 5.9.
Under the same setting of Theorem 5.1, Algorithm 1 leads to the following regret bounds over $N$ iterations:
$\sum_{n=1}^{N}\mathbb{E}\big{[}\bar{J}(\boldsymbol{\phi}^{*},\boldsymbol{0})-\bar{J}(\boldsymbol{\phi}_{n},\Gamma_{n})\big{]}\leq c+cN^{\frac{3}{4}}(\log N)^{\frac{p_{1}+p_{2}}{2}+1}(\log\log N),\quad\text{if }x_{0}\neq 0,$
$\sum_{n=1}^{N}\mathbb{E}\big{[}\bar{J}(\boldsymbol{\phi}^{*},\boldsymbol{0})-\bar{J}(\boldsymbol{\phi}_{n},\Gamma_{n})\big{]}\leq c+cN^{\frac{3}{4}}(\log N)^{\frac{p_{1}^{\prime}+p_{2}}{2}\vee(p_{1}^{\prime}+\frac{3}{2})}(\log\log N),\quad\text{if }x_{0}=0,$
where $c>0$ is a constant independent of $N$, and $p_{1},p_{1}^{\prime},p_{2}$ are the same constants given in Theorems 5.1 and 5.8.
Proof.
Following [1, Lemma B.7, B.8], we have
$\bar{J}(\boldsymbol{\phi},\Gamma)=f(a(\boldsymbol{\phi}))+(\sum_{j=1}^{m}D_{j}^{\top}\Gamma D_{j})g(a(\boldsymbol{\phi})),$ (54)
where
$a(\boldsymbol{\phi})=2A+2B^{\top}\boldsymbol{\phi}+\sum_{j=1}^{m}(C_{j}^{2}+2C_{j}D_{j}^{\top}\boldsymbol{\phi}+D_{j}^{\top}\boldsymbol{\phi}\boldsymbol{\phi}^{\top}D_{j}),$ (55)
and
$f(a)=\begin{cases}\frac{x_{0}^{2}(-H-QT)}{2}&\text{if }a=0,\\ \frac{1}{2a}(Q-e^{aT}Q-He^{aT}a)x_{0}^{2}&\text{if }a\neq 0,\end{cases}$ (56)
$g(a)=\begin{cases}\frac{T(-2H-QT)}{4}&\text{if }a=0,\\ \frac{1}{2a^{2}}(QTa+Q+Ha-e^{aT}Q-He^{aT}a)&\text{if }a\neq 0.\end{cases}$ (57)
Clearly, both $f$ and $g$ are continuously differentiable, non-increasing, and strictly negative everywhere.
Next, we present the proofs under $x_{0}\neq 0$ and $x_{0}=0$.
Case I: $x_{0}\neq 0.$ By (54), we have
$\bar{J}(\boldsymbol{\phi}^{*},0)-\bar{J}(\boldsymbol{\phi}_{n},\Gamma_{n})=$ $[f(a(\boldsymbol{\phi}^{*}))-f(a(\boldsymbol{\phi}_{n}))]-(\sum_{j=1}^{m}D_{j}^{\top}\Gamma_{n}D_{j})g(a(\boldsymbol{\phi}^{*}))$ (58)
$+(\sum_{j=1}^{m}D_{j}^{\top}\Gamma_{n}D_{j})[g(a(\boldsymbol{\phi}^{*}))-g(a(\boldsymbol{\phi}_{n}))].$
The goal is to establish bounds for the three terms on the right-hand side of (58), adapting the arguments of the proof of [1, Theorem 4.2] to the current setting. The first term, $f(a(\boldsymbol{\phi}^{*}))-f(a(\boldsymbol{\phi}_{n}))$, can be bounded exactly as in the proof of [1, Theorem 4.2], since it does not involve $\Gamma_{n}$. For the second term, $\left(\sum_{j}D_{j}^{\top}\Gamma_{n}D_{j}\right)g(a(\boldsymbol{\phi}^{*}))$, and the third term, $\left(\sum_{j}D_{j}^{\top}\Gamma_{n}D_{j}\right)[g(a(\boldsymbol{\phi}^{*}))-g(a(\boldsymbol{\phi}_{n}))]$, although $\Gamma_{n}$ is now stochastic rather than deterministic, the bounds can still be derived by utilizing part (b) of Theorem 5.1 and applying the Hölder and Cauchy–Schwarz inequalities. These modifications allow the original argument in the proof of [1, Theorem 4.2] to follow through.
Case II: $x_{0}=0$. It follows from (56) that $f(a(\boldsymbol{\phi}))=0$ for any $\boldsymbol{\phi}$. Hence
$\bar{J}(\boldsymbol{\phi}^{*},0)-\bar{J}(\boldsymbol{\phi}_{n},\Gamma_{n})=-(\sum_{j=1}^{m}D_{j}^{\top}\Gamma_{n}D_{j})g(\boldsymbol{\phi}^{*})+(\sum_{j=1}^{m}D_{j}^{\top}\Gamma_{n}D_{j})(g(\boldsymbol{\phi}^{*})-g(\boldsymbol{\phi}_{n})).$ (59)
The bound for the term $\left(\sum_{j}D_{j}^{\top}\Gamma_{n}D_{j}\right)g(a(\boldsymbol{\phi}^{*}))$ can be derived exactly as with the second term in Case I. Moreover, utilizing the MSE of $\boldsymbol{\phi}_{n}$ for $x_{0}=0$, we can also derive the bound for the other term, $(\sum_{j}D_{j}^{\top}\Gamma_{n}D_{j})(g(\boldsymbol{\phi}^{*})-g(\boldsymbol{\phi}_{n}))$, analogously to Case I.
Remark 5.10.
Discretization errors arising from the time step size $\Delta t_{n}$ are captured by the terms $\beta^{(\boldsymbol{\phi})}_{n}$ and $\beta^{(\Gamma)}_{n}$. Theorems 5.3 - 5.8 suggest that $\beta^{(\boldsymbol{\phi})}_{n}$ and $\beta^{(\Gamma)}_{n}$ should be set to decrease at the order of $n^{-\frac{3}{8}}$, which can be achieved by setting $\Delta t_{n}=T(n+1)^{-\frac{5}{8}}$. We omit the details here, but refer to [39, 25, 1] for discussions on time-discretization analyses.
6 Numerical Experiments
This section reports four numerical experiments to evaluate the performance of our proposed exploratory adaptive LQ-RL algorithm. The first one validates our theoretical findings by examining the convergence behaviors of $\boldsymbol{\phi}$ and $\Gamma$ and confirming the sublinear regret bound. The second one compares our model-free approach with a model-based benchmark ([25]) adapted to our setting that incorporates state- and control-dependent volatility. The third one evaluates the effect of exploration strategies by comparing our approach to a deterministic exploration schedule [1]. We examine scenarios where exploration parameters are either overly conservative or excessively aggressive, demonstrating the advantages of dynamically adjusting exploration based on observed data rather than relying on predefined schedules. The final experiment extends the third one by introducing random model parameters and random initial exploration parameter values, capturing real-world situations where no prior knowledge of the environment is available. This last experiment further demonstrates the effectiveness of the data-driven exploration mechanism in adapting to the environment.
For all the experiments, we set the control dimension and Brownian motion dimension to be $l=1$ and $m=1$. Under this setting, the LQ dynamics are simplified to
$\mathrm{d}x^{u}(t)=(Ax^{u}(t)+Bu(t))\mathrm{d}t+(Cx^{u}(t)+Du(t))\mathrm{d}W(t).$ (60)
Moreover, we set all the model parameters $A,B,C,D,Q,H,x_{0},$ and $T$ to be 1 and use a time step of $\Delta t=0.01$ for the first three experiments. See also Appendix A for the replicability of our experiments.
6.1 Numerical Validation of Theoretical Results
To validate the theoretical results on the convergence rates and regret bounds, we focus on the actor parameters $\Gamma$ and $\boldsymbol{\phi}$. We conduct 100 independent runs, each consisting of 100,000 iterations to capture long-term behavior. Convergence rates and regret are plotted in Figure 1 using a log-log scale, where the logarithm of the MSE (for convergence) or regret is plotted against the logarithm of the number of iterations.
(a) MSE of $\Gamma$
(b) MSE of $\boldsymbol{\phi}$
(c) Regret
Figure 1: Log-log plot of Algorithm 1.
Figures 1(a) and 1(b) demonstrate that the convergence rates for $\Gamma$ and $\boldsymbol{\phi}$ are $-0.51$ and $-0.52$, respectively, which closely match the theoretical results in Theorems 5.1 and 5.8 respectively. Additionally, Figure 1(c) shows a regret slope of $0.73$, aligning well with Theorem 5.9. Overall, these numerical results are consistent with the theoretical analysis, confirming the long-term effectiveness of our learning algorithm.
6.2 Model-Free vs. Model-Based
Under the same setting, we compare our model-free LQ-RL algorithm with data-driven exploration to the recently developed model-based method with a deterministic exploration schedule [25], which estimates parameters $A$ and $B$ in the drift term under the assumption of a constant volatility. To cover the settings with state- and control-dependent volatilities, [1] modified the algorithm in [25] by incorporating estimations of also $C$ and $D$, which we adopt in our experiment.
(a) MSE of $\boldsymbol{\phi}$ in model-based algorithm
(b) Regret of model-based algorithm
Figure 2: Log-log plot of model-based LQ-RL algorithm with fixed exploration schedule.
[…]
6.3 Adaptive vs. Fixed Explorations
Now, we compare two model-free continuous-time LQ-RL algorithms: ours with data-driven exploration (LQRL_Adaptive) and the one by [1] with fixed exploration schedules (LQRL_Fixed). Both algorithms follow the same fundamental framework, with the key distinction being how exploration is carried out. In Section 3.1, we outlined several drawbacks associated with deterministic exploration schedules. To illustrate them more concretely, we analyze two scenarios where a pre-determined exploration is either excessive or insufficient.
For comparison, we examine the trajectories of the actor exploration parameter $\Gamma$ and the cumulative regrets over iterations. [^4] The temperature parameter $\gamma$ is not compared, as it remains fixed in [1]. Both algorithms share exactly the same experimental setup, and we conduct 1,000 independent runs to observe the overall performances.
Excessive Exploration with a Near-Optimal Policy
We first consider a scenario where the policy parameter is initialized at $\boldsymbol{\phi}_{0}=-1.8$, which is close to its optimal value $\boldsymbol{\phi}^{*}=-2$. As the initial policy is already nearly optimal, exploration is not highly necessary. However, the initial exploration levels by both the critic and actor are set excessively high at $\gamma_{0}=\Gamma_{0}=20$.
(a) Trajectory of $\Gamma$
(b) Cumulative regret
Figure 3: Comparison of exploration level and cumulative regret under excessive initial exploration.
Figure 3(a) shows that while both algorithms start with an unnecessarily high level of $\Gamma$, LQRL_Adaptive rapidly adjusts $\Gamma$ downward, having effectively learned that exploration is less needed in this setting. By contrast, LQRL_Fixed follows a predetermined decay in $\Gamma$ and remains at a consistently higher level. As a result, Figure 3(b) demonstrates that LQRL_Adaptive achieves lower cumulative regret, indicating a more efficient learning process by dynamically adapting exploration to the circumstances.
Insufficient Exploration with a Poor Policy
We now examine an opposite scenario where $\boldsymbol{\phi}_{0}$ starts at 00, relatively far away from its optimal value $\boldsymbol{\phi}^{*}=-2$. The initial exploration parameters are set too low at $\gamma_{0}=\Gamma_{0}=0.02$, creating a situation where increased exploration is crucial for effective learning.
(a) Trajectory of $\Gamma$
(b) Cumulative regret
Figure 4: Comparison of exploration level and cumulative regret under insufficient initial exploration.
Figure 4(a) shows that LQRL_Adaptive rapidly increases $\Gamma$ from its initial low value of $0.02$ to nearly $1$, maintaining a higher exploration level than LQRL_Fixed throughout. By contrast, LQRL_Fixed follows a predetermined decaying schedule, reducing $\Gamma$ even further over iterations, which is counterproductive in this case where greater exploration is called for. As a result, Figure 4(b) demonstrates an even larger performance gap in cumulative regret compared to the previous scenario, further underscoring the benefits of adaptively adjusting exploration for learning.
6.4 Adaptive vs. Fixed Explorations: Randomized Model Parameters
In the final experiment, we evaluate the robustness of the comparison between LQRL_Adaptive and LQRL_Fixed in a setting where model parameters and initial exploration levels are randomly generated, mimicking real-world scenarios with no prior knowledge of the environment and no pre-tuning of exploration parameters. Specifically, for each simulation run, the model parameters $A,B,C,$ and $D$ are sampled from a uniform distribution $\mathcal{U}(-5,5)$, while the initial exploration levels of $\gamma_{0}$ and $\Gamma_{0}$ follow $\mathcal{U}(0,5)$. The initial policy parameter is set to be $\boldsymbol{\phi}_{0}=0$ as a neutral starting point. To ensure the reliability of the experiment, we conduct $10,000$ independent runs.
Figure 5: Comparison of regrets under randomized environments.
Figure 5 demonstrates that LQRL_Adaptive significantly outperforms LQRL_Fixed, with the performance gap widening as the number of iterations increases.
7 Conclusions
This paper is a continuation of [1], aiming to develop a model-free continuous-time RL framework for a class of stochastic LQ problems with state- and control-dependent volatilities. A key contribution, compared to [1], is the introduction of a data-driven exploration mechanism that adaptively adjusts both the temperature parameter controlled by the critic and the stochastic policy variance managed by the actor. This approach overcomes the limitations of fixed exploration schedules, which often require extensive tuning and incur unnecessary exploration costs. While the theoretically proved sublinear regret bound of $O(N^{\frac{3}{4}})$ is the same as that in [1], numerical experiments confirm that our adaptive exploration strategy significantly improves learning efficiency compared to [1].
Research on model-free continuous-time RL is still in the very early innings. To our best knowledge [1, 40] are the only works on regret analysis in the realm, largely because the problem is extremely challenging. We focus on the same LQ problem in [1] because this is the problem that we are able to solve at the moment, one nevertheless that may serve as a starting point for tackling more complex problems. Meanwhile, it remains an interesting open question whether adaptive exploration strategies also theoretically improve the regret bound, which is observed from our numerical experiments. Finally, our framework relies on entropy regularization and policy variance for exploration, which may be enhanced by goal-based constraints/rewards to improve sampling efficiency. All these provide promising avenues for future research, driving further advancements in data-driven exploration and continuous-time RL.
Appendix A More Details for Numerical Experiments
To ensure full replicability, we set the random seed from 1 to 100 for each independent run in both LQRL_Adaptive and the model-based benchmark in the first and second experiments. For the third experiment comparing LQRL_Adaptive and LQRL_Fixed, we set the random seed from 1 to 1,000 for each independent run in both scenarios. For the fourth experiment under the randomized environment, we set the random seed from 1 to 10,000.
The setup for our first experiment using the model-free algorithm with data-driven exploration (LQRL_Adaptive), Algorithm 1, is as follows. The initial policy mean parameter is set to $\boldsymbol{\phi}_{0}=-1.1$, with the actor and critic exploration levels initialized at $\Gamma_{0}=0.5$ and $\gamma_{0}=2$, respectively. The learning rates are specified by $a^{(\boldsymbol{\phi})}_{n}=\frac{0.05}{(n+1)^{3/4}}$ and $a^{(\Gamma)}_{n}=\frac{1}{(n+1)^{3/4}}$. For computational efficiency, the projection sets are defined as $[-2.25,-1.1]$ for $\boldsymbol{\phi}_{n}$ and $[0,1]$ for $\Gamma_{n}$. Although the projections were originally specified for theoretical convergence guarantees, they are slightly modified in our experiments to accelerate convergence. The projection sets for $\boldsymbol{\theta}$ and $\gamma$ remain unbounded. The deterministic sequence $b_{n}$ is chosen as $b_{n}=20\max(1,(n+1)^{1/4})$. Additionally, we define ${k}_{1}(t;\boldsymbol{\theta})=1$ and ${k}_{3}(t;\boldsymbol{\theta},\gamma)=0$, satisfying the conditions in (11), noting that our results do not depend on the explicit form of the value function.
The model-based benchmark is adapted from [25] and extended by [1, Algorithm A.1] to accommodate state- and control-dependent volatility in our setting. The initial parameter estimations are set to be $A=B=C=D=10$, resulting in the same initial $\boldsymbol{\phi}_{0}=-1.1$ as in LQRL_Adaptive. The initial exploration level $\Gamma_{0}=0.5$ is also matched. The deterministic exploration schedule follows $\Gamma_{n}=0.5(n+1)^{-1}$, consistent with [1]. Finally, the projection set for $\boldsymbol{\phi}_{n}$ remains $[-2.25,-1.1]$, aligning with the setup used in LQRL_Adaptive.
For all the plots in the first and second experiments, for both our adaptive model-free algorithm and model-based benchmarks, regression lines are fitted using data from iterations 5,000 to 100,000. This avoids the influence of early-stage learning noise. In all the regret plots, we use the median and exclude extreme values. [^5] Similar conclusions hold when using the mean instead.
In our third experiment, comparing LQRL_Adaptive and LQRL_Fixed, most settings remain identical to the first experiment, except that the bounds $c^{(\boldsymbol{\phi})}_{n}$ and $c^{(\Gamma)}_{n}$ are set to be 20 to accommodate the scenarios under consideration. For the fourth experiment, to account for potentially large values of $\boldsymbol{\phi}$ and $\Gamma$ due to random model parameters, we further increase the bounds $c^{(\boldsymbol{\phi})}_{n}$ and $c^{(\Gamma)}_{n}$ to be 100.
Acknowledgment
The authors are supported by the Nie Center for Intelligent Asset Management at Columbia University. Their work is also part of a Columbia- CityU/HK collaborative project that is supported by the InnoHK Initiative, The Government of the HKSAR, and the AIFT Lab.